OpenStudy (anonymous):

how to simplify ∛16a^4b^3

OpenStudy (anonymous):

=2ab∛2a

OpenStudy (anonymous):

you have to use a calculator for this

OpenStudy (anonymous):

I'm confused on how to simplify the 16

OpenStudy (anonymous):

you plug the 3radical sign thingy and 16 into your calculator, if you have a calculator that can do that, therefor it will give you the correct answer

OpenStudy (anonymous):

To take a cube root you need 3 repeated factors under the sign. In this case: \[\sqrt[3]{16a^4b^3}=\sqrt[3]{2\times2\times2\times2\times a \times a \times a \times a \times b \times b \times b}\] because there are 4 2's you can pull a 2 (3 2's under the sign) out leaving 1 true also for the a's you can pull out a b leaving none this becomes \[2ab \sqrt[3]{2a}\]